# On elliptic and quasiregularly elliptic manifolds

### v1

@@ 1. Introduction
Language barriers between different mathematical subfields often lead to the same ideas being developed independently in different contexts. In this paper, we draw renewed attention to the parallels between elliptic and quasiregularly elliptic manifolds. Both properties are defined by the existence of certain maps from Euclidean spaces; the two types of maps naturally lend themselves to different tools. Although a link between the two notions was already suggested by Gromov in [13, §2.E and 6.D], similar topological obstructions for ellipticity and quasiregular ellipticity have since been developed in parallel, among others by the two authors. The similarity between both the statements and the proofs in the two cases supports Gromov’s intuition.
@@ 1.1. Definitions and their motivations
Gromov [13, 2.41] suggested calling a manifold $M$ elliptic [^1] This terminology is inspired by, but should not be confused with, the rather distinct notion of elliptic spaces from rational homotopy theory. We sometimes say “Gromov-elliptic” to emphasize this difference. if there is a Lipschitz map $f:\mathbb{R}^{n}\to M$ of positive asymptotic degree. One might as well demand that $f$ be $1$-Lipschitz, in which case the map can be thought of as efficiently wrapping $M$ with an infinite sheet of wrapping paper; to be more concise, we will call such a map a wrapping map.
Recently [6, 5], ellipticity has been shown to be crucial to understanding Lipschitz maps to $M$ from other manifolds. For example, given a nullhomotopic $L$-Lipschitz map $N\to M$, is there a $C(L+1)$-Lipschitz nullhomotopy, for some constant $C=C(M,N)$? This turns out to be true if $M$ is elliptic, simply connected, and formal, a condition from rational homotopy theory, and false if $M$ is simply connected and formal but not elliptic. Spaces satisfying all three conditions are called scalable. More general conjectures of Gromov about how the rational homotopy class of a map is constrained by its Lipschitz constant [13, Ch. 7] are likewise confirmed for scalable spaces and false for spaces which are simply connected and formal but not scalable.
@@
Quasiregular maps are discrete (in the sense that the preimage of every point is a discrete set) and open due to a theorem of Reshetnyak [33]. They should therefore be thought of as generalized branched coverings with local geometric control. A manifold is said to be quasiregularly (or QR) elliptic, a term introduced by Bonk and Heinonen [7] following Gromov’s template, if it admits a quasiregular map from $\mathbb{R}^{n}$.
@@
Following Le Donne and Pankka [20], we say that a Riemannian $n$-manifold $M$ is BLD-elliptic if it admits a BLD map from $\mathbb{R}^{n}$. Since BLD maps $\mathbb{R}^{n}\to M$ are both quasiregular and wrapping maps, every BLD-elliptic manifold is both elliptic and quasiregularly elliptic. On the other hand, Pankka has suggested that every quasiregularly elliptic manifold may be BLD-elliptic. This would explain many of the seeming coincidences observed in this paper.
@@ 1.2. Obstructions to ellipticity and quasiregular ellipticity
Theorem 1.1 (Varopoulos [36, Theorem X.5.1], Gromov [13, Theorem 2.42]).
@@
Gromov gives several other obstructions to ellipticity in [13, §2.E], most of which can be derived from Theorem 1.1 and the following other result:
Theorem 1.2 (Berdnikov–Guth–Manin [5, Theorem 2.3], Heikkilä–Pankka [15]).
@@
In the elliptic case, this is a theorem of Berdnikov, Guth and Manin, who also show a quantitative refinement [5, Theorem 2.3]. The quasiregular result is a refinement of Prywes’ main theorem in [31]; after we began our work an independent proof was posted by Heikkilä and Pankka [15]. We give relatively concise and similar proofs of both statements by combining ideas from [6] and [31]. In both cases, the main idea is to take pullbacks of arbitrary differential forms generating the cohomology ring and take a scaling limit. While the forms may not obey the same relations as the cohomology classes, their scaling limits do, giving a ring homomorphism from $H^{*}(M;\mathbb{R})$ to an algebra of forms on $\mathbb{R}^{n}$. Evaluating this ring homomorphism at a point gives the desired result.
We give a number of corollaries of this result, the most important of which is the following, answering a question of Gromov [13, Question 2.44]:
@@
The analogous result for BLD-elliptic manifolds was obtained using different methods by Le Donne and Pankka [20].
Most known obstructions to ellipticity and quasiregular ellipticity follow from Theorem 1.1, Theorem 1.2, and their corollaries. The main exception is the following:
Theorem 1.4 (Gromov [13, Theorem 2.45], Luisto and Pankka [21]).
Suppose that $M$ is an elliptic or quasiregularly elliptic closed $n$-manifold whose fundamental group is virtually $\mathbb{Z}^{n}$. Then $M$ is aspherical (and therefore finitely covered by a torus).
This theorem is particularly interesting because unlike Theorem 1.2, it gives torsion information. For example, it shows that $T^{n}\mathbin{\#}\Sigma$ is not elliptic or quasiregularly elliptic, where $\Sigma$ is a nontrivial simply connected rational homology sphere. [^2] Such a $\Sigma$ can be constructed as follows. By attaching a $3$-cell to $S^{2}$ along a map of degree $2$, one creates a CW complex $X$ whose only nonzero homology group is $H_{2}(X)\cong\mathbb{Z}/2\mathbb{Z}$. By embedding $X$ in $\mathbb{R}^{7}$ and taking the boundary of a neighborhood of $X$, we obtain a $6$-manifold whose homology is the same as that of $X$ in low dimensions. By Poincaré duality, this manifold is a rational homology sphere. The projection map from $T^{n}\mathbin{\#}\Sigma$ to $T^{n}$ is an isomorphism on rational cohomology.
In contrast, for simply connected spaces, we give the following new result:
@@
A nilpotent space is a space whose fundamental group is nilpotent and acts nilpotently on its higher homotopy groups. In particular, the higher homotopy groups of a nilpotent space are all finitely generated. On the other hand, for the manifold $T^{n}\mathbin{\#}\Sigma$ discussed above, if $k>1$ is minimal such that $H_{k}(\Sigma)\neq 0$, then
$\pi_{k}(T^{n}\mathbin{\#}\Sigma)\cong\pi_{k}(\widetilde{T^{n}\mathbin{\#}\Sigma})\cong H_{k}(\widetilde{T^{n}\mathbin{\#}\Sigma})\cong\mathbb{Z}^{n}[H_{k}(\Sigma)].$
@@
Proof.
Suppose that the $n$-manifold $M$ is elliptic, and let $f:\mathbb{R}^{n}\to M$ be a wrapping map. Suppose that $N$ has the same rational homotopy type as $M$. We use Sullivan’s model of rational homotopy theory (see [35, 11, 9]). The manifolds $M$ and $N$ have the same Sullivan minimal algebra $\mathcal{A}$ and minimal models $m_{M}:\mathcal{A}\to\Omega^{*}M$ and $m_{N}:\mathcal{A}\to\Omega^{*}N$. By the shadowing principle [24, Theorem 4.1], [^3] This is proved in [24] only for simply connected target spaces. The generalization to nilpotent spaces in contained in forthcoming work of Kyle Hansen. the DGA homomorphism $f^{*}m_{M}:\mathcal{A}\to\Omega^{*}\mathbb{R}^{n}$ extends to a DGA homomorphism $\Phi:\mathcal{A}\to\Omega^{*}\mathbb{R}^{n}\times[0,1]$ such that
  1. (i)
    The forms in the image of $\Phi$ have bounded operator norm, and
  2. (ii)
    $\Phi|_{t=1}=g^{*}m_{N}$ for some Lipschitz map $g:\mathbb{R}^{n}\to N$.
In particular, since
it follows that $g$ is also a wrapping map. ∎
@@ 1.3. Construction of elliptic and quasiregularly elliptic manifolds
A general construction is given in [5] for formal manifolds. A simply connected space is formal in the sense of Sullivan [35, §12] if, informally speaking, its rational homotopy type is the simplest one with its rational cohomology ring. An equivalent informal condition is that there are no nontrivial higher-order rational cohomology operations: the only relations between cohomology classes are given by the cup product. For finite complexes, another useful equivalent condition is as follows: a simply connected finite CW complex $X$ is formal if and only if there is an integer $p>1$ and a map $f_{p}:X\to X$ whose induced map on $H^{k}(X;\mathbb{Q})$, for every $k$, is multiplication by $p^{k}$.
Examples of formal manifolds include all simply connected symmetric spaces, such as spheres, complex projective spaces, and Grassmannians. Products and connected sums of formal manifolds are also formal. The simplest and lowest-dimensional example of a non-formal simply connected manifold is the total space $M$ of a fiber bundle $S^{3}\to M\to S^{2}\times S^{2}$ obtained by pulling back the Hopf fibration $S^{3}\to S^{7}\to S^{4}$ along the degree $1$ map $S^{2}\times S^{2}\to S^{4}$ [9, p. 94].
Theorem 1.7 ([5, Theorem C]).
@@
We note that if there is an embedding $H^{*}(M;\mathbb{Q})\hookrightarrow\Lambda^{*}\mathbb{Q}^{n}$, then there is a map of positive degree from $T^{n}$ to $M$, and therefore $M$ is trivially elliptic. However, a real embedding as above need not be rationalizable; a 9-dimensional example manifold $M$ is given in [5, Remark 5.2]. A wrapping map of this $M$ must therefore be in some sense aperiodic (but may be almost periodic).
@@
  • In dimension $3$, only $S^{3}$, $S^{2}\times S^{1}$, $T^{3}$, and their finite quotients are quasiregularly elliptic due to work of Jormakka [19] and the geometrization theorem.
  • In dimension $4$, Heikkilä and Pankka [15] give a classification up to homeomorphism in the simply connected case: all those manifolds which are not excluded by Theorem 1.2 are homeomorphic to quasiregularly elliptic ones by work of Piergallini and Zuddas [30] because they admit a branched covering by $T^{4}$. Specifically, these are
    #$\#^{r}(S^{2}\times S^{2}),\qquad\#^{s}\mathbb{C}P^{2}\#^{t}\overline{\mathbb{C}P^{2}},\qquad 0\leq r,s,t\leq 3,$
@@
    Note that this is not a complete classification of which $4$-manifolds are quasiregularly elliptic, since quasiregular ellipticity is only known to be invariant up to PL homeomorphism. Some of the manifolds above are known to admit exotic smooth (and therefore PL) structures [2, 3]; it is not known whether any such exotic $4$-manifolds are quasiregularly elliptic.
In higher dimensions, it is known that $\mathbb{C}P^{n}$ is quasiregularly elliptic, as shown in Prywes’s dissertation [32, Prop. 4.1.3].
@@ 1.4. Open manifolds
Thus far we have discussed Gromov- and quasiregular ellipticity for closed manifolds. This restriction is natural because it makes the question topological: for closed manifolds, ellipticity is a homotopy invariant, [^4] This follows from the fact that any homotopy equivalence between compact Riemannian manifolds can be made Lipschitz. and quasiregular ellipticity is invariant at least under PL homeomorphism [15, Proposition B.1]. On the other hand, for open manifolds both properties depend on the metric, and in different ways, as we discuss further in §4 and 5.
In prior works on quasiregular ellipticity such as [28, 29], an open manifold is said to be quasiregularly elliptic if it admits a quasiregular map from $\mathbb{R}^{n}$ with respect to some Riemannian metric. This forces quasiregular ellipticity to be a topological property. Even with this definition, Theorem 1.1 extends to the open manifold case:
Theorem 1.8 (Pankka–Rajala [28, Theorem 1.3]).
@@
The hypothesis holds, for example, for open subsets of closed manifolds, as well as for manifolds of pinched negative curvature and nonpositively curved locally symmetric spaces. The easiest non-examples are infinite strings of surfaces joined by thinner and thinner necks. In future work it will be shown that it is equivalent to $M$ having a nonzero Cheeger constant.
Even then, there are manifolds of dimension $n\geq 3$ which have an elliptic metric satisfying this hypothesis, but no quasiregularly elliptic metric. One such example is $S^{n}\setminus X$ where $X$ is a countable subset of the equatorial $S^{n-2}$. We show in Theorem 5.1 that $S^{n}\setminus S^{n-2}$ (with the round metric) is elliptic, and therefore so is $S^{n}\setminus X$. On the other hand, a generalization of the Rickman–Picard theorem [34] by Holopainen and Rickman [18, Theorem 3.1] shows that $S^{n}\setminus X$ cannot be quasiregularly elliptic in the topological sense above.

### v3

@@ 1. Introduction
Language barriers between mathematical subfields often lead to the same ideas being developed independently in different contexts. In this paper, we draw renewed attention to the parallels between elliptic and quasiregularly elliptic manifolds. Both properties are defined by the existence of certain maps from Euclidean spaces; the two types of maps naturally lend themselves to different tools. Although a link between the two notions was already suggested by Gromov in [GrMS, §2.E and 6.D], similar topological obstructions for ellipticity and quasiregular ellipticity have since been developed in parallel, among others by the two authors. The similarity between both the statements and the proofs in the two cases supports Gromov’s intuition.
@@ 1.1. Definitions and their motivations
Gromov [GrMS, 2.41] suggested calling a manifold $M$ elliptic [^1] This terminology is inspired by, but should not be confused with, the rather distinct notion of elliptic spaces from rational homotopy theory. We sometimes say “Gromov-elliptic” to emphasize this difference. if there is a Lipschitz map $f:\mathbb{R}^{n}\to M$ of positive asymptotic degree. One might as well demand that $f$ be $1$-Lipschitz, in which case the map can be thought of as efficiently wrapping $M$ with an infinite sheet of wrapping paper; to be more concise, we will call such a map a wrapping map.
Recently [BeM, BeGM], ellipticity has been shown to be crucial to understanding Lipschitz maps to $M$ from other manifolds. For example, given a nullhomotopic $L$-Lipschitz map $N\to M$, is there a $C(L+1)$-Lipschitz nullhomotopy, for some constant $C=C(M,N)$? This turns out to be true if $M$ is elliptic, simply connected, and formal, a condition from rational homotopy theory, and false if $M$ is simply connected and formal but not elliptic. Spaces satisfying all three conditions are called scalable. More general conjectures of Gromov about how the rational homotopy class of a map is constrained by its Lipschitz constant [GrMS, Ch. 7] are likewise confirmed for scalable spaces and false for spaces which are simply connected and formal but not scalable.
@@
Quasiregular maps are discrete (in the sense that the preimage of every point is a discrete set) and open due to a theorem of Reshetnyak [Reshet]. They should therefore be thought of as generalized branched coverings with local geometric control. A manifold is said to be quasiregularly (or QR) elliptic, a term introduced by Bonk and Heinonen [BH] following Gromov’s template, if it admits a quasiregular map from $\mathbb{R}^{n}$.
@@
Following Le Donne and Pankka [LDPa], we say that a Riemannian $n$-manifold $M$ is BLD-elliptic if it admits a BLD map from $\mathbb{R}^{n}$. Since BLD maps $\mathbb{R}^{n}\to M$ are both quasiregular and wrapping maps, every BLD-elliptic manifold is both elliptic and quasiregularly elliptic. On the other hand, Pankka has suggested that every quasiregularly elliptic manifold may be BLD-elliptic. This would explain many of the seeming coincidences observed in this paper.
@@ 1.2. Obstructions to ellipticity and quasiregular ellipticity
Theorem 1.1 (Varopoulos [vsc, Theorem X.5.1], Gromov [GrMS, Theorem 2.42]).
@@
The other main obstruction to ellipticity is the following:
Theorem 1.2 (Berdnikov–Guth–Manin [BeGM, Theorem 2.3], Heikkilä–Pankka [HP]).
@@
In the elliptic case, this is a theorem of Berdnikov, Guth and Manin, who also show a quantitative refinement [BeGM, Theorem 2.3]. The quasiregular result is a refinement of Prywes’ main theorem in [Pry] first shown by Heikkilä and Pankka [HP]. We give relatively concise and similar proofs of both statements by combining ideas from [BeM] and [Pry]. In both cases, the main idea is to take pullbacks of arbitrary differential forms generating the cohomology ring and take a scaling limit. While the forms may not obey the same relations as the cohomology classes, their scaling limits do, giving a ring homomorphism from $H^{*}(M;\mathbb{R})$ to an algebra of forms on $\mathbb{R}^{n}$. Evaluating this ring homomorphism at a point gives the desired result.
Most known obstructions to ellipticity and quasiregular ellipticity, including some obstructions to ellipticity derived by Gromov in [GrMS, §2.E], follow from Theorem 1.1 and Theorem 1.2. The most far-reaching such corollary is the following, which answers a question of Gromov [GrMS, Question 2.44]:
@@
The analogous result for BLD-elliptic manifolds was obtained using different methods by Le Donne and Pankka [LDPa].
We give several further corollaries in §3, of which the easiest to state are the following. Let $M$ be an elliptic or quasiregularly elliptic $n$-manifold. Then:
  • The fundamental group $\pi_{1}(M)$ cannot have rank n1n-1.
  • If $\pi_{1}(M)$ is infinite, then the Euler characteristic $\chi(M)=0$.
The latter is also shown in [PG]; the authors use a different method (studying the growth of Betti numbers of finite covers), but their result is likewise a purely topological corollary of Theorem 1.1 and known results about the cohomology ring.
In fact, as far as we know, the only obstruction to ellipticity or quasiregular ellipticity of closed manifolds which does not follow from Theorems 1.1 and 1.2 is the following:
Theorem 1.4 (Gromov [GrMS, Theorem 2.45]).
Suppose that $M$ is an elliptic closed $n$-manifold whose fundamental group is virtually $\mathbb{Z}^{n}$. Then $M$ is aspherical (and therefore finitely covered by a manifold homeomorphic to the torus).
The corresponding statement for quasiregularly elliptic manifolds is believed to be true, and an argument was given by Luisto and Pankka [LuPa]. However, after the first version of this article was posted, Pankka noticed that their argument only showed that $M$ must have a rationally acyclic universal cover.
Theorem 1.4 is particularly interesting because unlike Theorem 1.2, it gives torsion information. For example, it shows that $T^{n}\mathbin{\#}\Sigma$ is not elliptic, where $\Sigma$ is a nontrivial simply connected rational homology sphere. [^2] Such a $\Sigma$ can be constructed as follows. By attaching a $3$-cell to $S^{2}$ along a map of degree $2$, one creates a CW complex $X$ whose only nonzero homology group is $H_{2}(X)\cong\mathbb{Z}/2\mathbb{Z}$. By embedding $X$ in $\mathbb{R}^{7}$ and taking the boundary of a neighborhood of $X$, we obtain a $6$-manifold whose homology is the same as that of $X$ in low dimensions. By Poincaré duality, this manifold is a rational homology sphere. The projection map from $T^{n}\mathbin{\#}\Sigma$ to $T^{n}$ is an isomorphism on rational cohomology.
In contrast, for a large class of spaces, we prove the following result:
@@
A nilpotent space is a space whose fundamental group is nilpotent and acts nilpotently on its higher homotopy groups. A prototypical example would be the mapping torus of a self-map of $(S^{n})^{r}$ which induces a unipotent transformation on $H_{n}$.
In particular, the higher homotopy groups of a nilpotent space are all finitely generated. On the other hand, for the manifold $T^{n}\mathbin{\#}\Sigma$ discussed above, if $k>1$ is minimal such that $H_{k}(\Sigma)\neq 0$, then
$\pi_{k}(T^{n}\mathbin{\#}\Sigma)\cong\pi_{k}(\widetilde{T^{n}\mathbin{\#}\Sigma})\cong H_{k}(\widetilde{T^{n}\mathbin{\#}\Sigma})\cong\mathbb{Z}^{n}[H_{k}(\Sigma)],$
so it is not nilpotent. In fact, though, the key ingredient of Theorem 1.4 seems to be that the fundamental group has full rank, since other similar non-nilpotent examples are elliptic:
Theorem 1.6.
Let $X=(T^{n-k}\times S^{k})\mathbin{\#}\Sigma$, where $2\leq k\leq n-1$ and $\Sigma$ is a simply connected rational homology sphere. Then $X$ is elliptic.
@@ 1.3. Construction of elliptic and quasiregularly elliptic manifolds
A general construction is given in [BeGM] for formal manifolds. A simply connected space is formal in the sense of Sullivan [SulLong, §12] if, informally speaking, its rational homotopy type is the simplest one with its rational cohomology ring. An equivalent informal condition is that there are no nontrivial higher-order rational cohomology operations: the only relations between cohomology classes are given by the cup product. For finite complexes, another useful equivalent condition is as follows: a simply connected finite CW complex $X$ is formal if and only if there is an integer $p>1$ and a map $f_{p}:X\to X$ whose induced map on $H^{k}(X;\mathbb{Q})$, for every $k$, is multiplication by $p^{k}$.
Examples of formal manifolds include all simply connected symmetric spaces, such as spheres, complex projective spaces, and Grassmannians. Products and connected sums of formal manifolds are also formal. The simplest and lowest-dimensional example of a non-formal simply connected manifold is the total space $M$ of an $S^{3}$-bundle over $S^{2}\times S^{2}$ [FOT, p. 94], further discussed in §1.7.
Theorem 1.8 ([BeGM, Theorem C]).
@@
We note that if there is an embedding $H^{*}(M;\mathbb{Q})\hookrightarrow\Lambda^{*}\mathbb{Q}^{n}$, then there is a map of positive degree from $T^{n}$ to $M$, and therefore $M$ is trivially elliptic. However, a real embedding as above need not be rationalizable; a 9-dimensional example manifold $M$ is given in [BeGM, Remark 5.2] and we review this example in §1.7. A wrapping map of this $M$ must therefore be in some sense aperiodic (but may be almost periodic).
@@
  • In dimension $3$, only $S^{3}$, $S^{2}\times S^{1}$, $T^{3}$, and their finite quotients are quasiregularly elliptic due to work of Jormakka [Jor] and the geometrization theorem.
  • In dimension $4$, Heikkilä and Pankka [HP] give a classification up to homeomorphism in the simply connected case: all those manifolds which are not excluded by Theorem 1.2 are homeomorphic to quasiregularly elliptic ones by work of Piergallini and Zuddas [PZ] because they admit a branched covering by $T^{4}$. Specifically, these are
    $^{r}(S^{2}\times S^{2}),\qquad\#^{s}\mathbb{C}P^{2}\#^{t}\overline{\mathbb{C}P^{2}},\qquad 0\leq r,s,t\leq 3,$
@@
    Note that this is not a complete classification of which $4$-manifolds are quasiregularly elliptic, since quasiregular ellipticity is only known to be invariant up to PL homeomorphism. Some of the manifolds above are known to admit exotic smooth (and therefore PL) structures [AP1, AP2]; it is not known whether any such exotic $4$-manifolds are quasiregularly elliptic.
In higher dimensions, it is known that $\mathbb{C}P^{n}$ is quasiregularly elliptic, as shown in Prywes’s dissertation [PryTh, Prop. 4.1.3].
@@ 1.4. Open manifolds
Thus far we have discussed Gromov- and quasiregular ellipticity for closed manifolds. This restriction is natural because it makes the question topological: for closed manifolds, ellipticity is a homotopy invariant, [^3] This follows from the fact that any homotopy equivalence between compact Riemannian manifolds can be made Lipschitz. and quasiregular ellipticity is invariant at least under PL homeomorphism [HP, Proposition B.1]. On the other hand, for open manifolds both properties depend on the metric, and in different ways, as we discuss further in §5 and 6.
In prior works on quasiregular ellipticity such as [PaRa, PaRaWu], an open manifold is said to be quasiregularly elliptic if it admits a quasiregular map from $\mathbb{R}^{n}$ with respect to some Riemannian metric. This forces quasiregular ellipticity to be a topological property. Even with this definition, Theorem 1.1 extends to the open manifold case:
Theorem 1.9 (Pankka–Rajala [PaRa, Theorem 1.3]).
@@
The hypothesis implies that any Lipschitz map from $\mathbb{R}^{n}$ equidistributes, that is, at large scales it “visits” any neighborhood roughly equally often—something which is always true of quasiregular maps. It holds, for example, for open subsets of closed manifolds, as well as for manifolds of pinched negative curvature and nonpositively curved locally symmetric spaces. The easiest non-examples are infinite strings of surfaces joined by thinner and thinner necks. In future work it will be shown that it is equivalent to $M$ having a nonzero Cheeger constant.
Even then, there are manifolds of dimension $n\geq 3$ which have an elliptic metric satisfying this hypothesis, but no quasiregularly elliptic metric. One such example is $S^{n}\setminus X$ where $X$ is a countable subset of the equatorial $S^{n-2}$. We show in Theorem 6.1 that $S^{n}\setminus S^{n-2}$ (with the round metric) is elliptic, and therefore so is $S^{n}\setminus X$. On the other hand, a generalization of the Rickman–Picard theorem [Rick] by Holopainen and Rickman [HoR2, Theorem 3.1] shows that $S^{n}\setminus X$ cannot be quasiregularly elliptic in the topological sense above.